OpenStudy (anonymous):

MEDAL FOR HELP! SWAG BELOW! A school fundraiser is selling a custom design t-shirt. Their revenues are modeled by the quadratic equation, R = 0.03t + 550, where R is revenue in dollars for the sale of t number of t-shirts. The expenses, E, for printing and selling t-shirts are given by the linear equation E = 250 + 3.50t If 75 students have already paid for a shirt, approximately how many more do they need to sell in order to break even A 25 t-shirts B 57 t shirts C 102 t-shirts D 177 t-shirts

OpenStudy (anonymous):

@dan815 @whpalmer4

OpenStudy (anonymous):

The answer is C, i know that but I want to know how. Can someone help explain :)

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

solve \[0.03t + 550=250 + 3.50t \] then subtract \(75\) from the answer

OpenStudy (anonymous):

its supposed to be .03t^2 my bad

OpenStudy (whpalmer4):

Okay, so the corrected revenue function is apparently\[R = 0.03t^2 + 550\]whereas expenses are given by \[E = 250 + 3.50t\] The break-even point is where revenues equal expenses: \[R=E\]\[0.03t^2+550=250+3.50t\] Solve that for \(t\) and subtract \(75\) to account for the \(75\) shirts already sold. Problem I have with all of this is that your equations must not be copied correctly, as those two functions are never equal for real values of \(t\). Solving @satellite73's version of the equations doesn't get you a working answer either. I'd appreciate seeing the correct problem, out of curiosity if nothing else!